# Exploration of morphological coherence in open clusters with “core-shell” structure

### v1

@@ 3.1 Distribution of morphological coherence of sample clusters
We obtained all parameters including the MD and ER of 132 sample clusters through ellipsoid fitting and expressions above defined, which are compiled a machine-readable and seen in Appendix 1. Fig. 4 displays the distributions of the MD with its error (left panel) as well as the ER with its error (right panel). We found that the errors of the MD of most sample clusters are less than 0.1, with most ER errors less than 0.2. Meanwhile, both errors, expressed as $\sigma_{d}$ and $\sigma_{e_{core}/e_{outer}}$, respectively, present relatively apparent positive correlations with the MD and ER. The main reason is that the fitting error is larger for the sample clusters with large MD or large ER.
@@
To get a preliminary view of the morphological coherence distribution of the sample clusters, we plotted the distribution between the MD and the ER of the sample clusters, as shown in Fig. 5. Each colored circle represents one sample cluster, with its color representative of the number of the sample clusters’ members. The gray bars (horizontal and vertical directions) cross it are the error bars of MD and ER, respectively. The horizontal dashed line in the Figure marks the position of ER equal to 1 where the ellipticity of the cluster’s internal morphology is consistent with the ellipticity of its external morphology. The picture shows an increasing tendency of the ER ($e_{core}$/$e_{outer}$) values as a function of the MD ($d$), which implies that sample clusters with more elliptical external morphology than internal morphology typically present smaller morphological dislocations (MD), see dark red filled circles in Fig. 5, and thus have greater morphological stability. Meanwhile, it is apparent that these clusters have a larger number of member stars. Therefore, we concluded that the number of sample clusters’ members plays a crucial role in keeping their morphological stability.
@@
Figure 5: The distribution between the MD of the sample clusters and their ER. Each colored circle represents one sample cluster, with its color being coded according to the logarithmic value of the number of member stars of the sample clusters. The horizontal and vertical gray bars indicate the errors in the MD of the clusters and their ER, respectively. The dashed line denotes the position of the smallest morphological difference between the clusters’ inner and outer regions.
@@
In addition, we found in Fig. 5 a sample cluster (HSC 2986) away from the distribution of most sample clusters. It is a special object for our sample population, because it has a relatively small number of members and the largest MD among all samples, although the error of the cluster’s MD is large. To examine the cluster further, we drew its 3D spatial distribution, as shown in Fig. 6. The left panel of the Figure presents the 3D spatial distributions and the fitted ellipsoids of the internal and external structures of the HSC 2986. Its internal and external structures were divided based on the radius of the half number of cluster members. In this panel, the black dashed line indicates the position of the MD of HSC 2986, with its length representing unnormalized MD, see Eq. 1. In order to verify whether the cluster really has a large unnormalized MD, we adopted the Gaussian Mixture Model (GMM) used by Tarricq et al. (2022) to distinguish among open clusters’ core region, tidal tail, and corona in two dimensions, with 1000 iterations to divide the members’ distribution of HSC 2986 into two components, as shown in the right panel of Fig. 6. The two components are purple and black dots in this picture. We then evaluated the density centers (purple pentagram and black pentagram) in the two parts of the cluster separately and finally derived its unnormalized MD, see the red dashed line in the right panel of Fig. 6.
@@
Figure 6: Left: 3D spatial structure of HSC 2986 and its ellipsoidal curves. The black dashed line indicates the unnormalized MD of HSC 2986. Other symbols and curves are the same as those in Fig. 2. Right: the 3D spatial structure of HSC 2986. The purple and black dots are divided by the Gaussian Mixture Model (GMM) method, with the red dashed line meaning the unnormalized MD of HSC 2986. The purple and black pentagrams are the maximum points of the density of the purple and black scatter distribution.
@@
By comparing the unnormalized MDs of HSC 2986 obtained, we can determine that HSC 2986 has a large unnormalized MD, which implies that the cluster is weakly morphologically stable. Moreover, we can see that the whole members’ distribution of HSC 2986 does not show a distinct core, presenting a diffuse spatial arrangement. Therefore, we speculated that the cluster may be susceptible to disruption by external forces, regardless of the cluster’s evolutionary state. However, it is also possible that its core and outer structures are two separate groups due to the large unnormalized MD, if HSC 2986 is misclassified as an open cluster.
@@
3.2 Correlations of the morphological coherence of sample clusters with number of members, and age
@@
Figure 7 displays the distributions between the morphological coherence (MD and ER) and the number of the clusters’ members as well as their ages, with their median filter curves. It is apparent, as seen in the (a) subplot of Fig. 7, that as the number of clusters’ members increases, both their MD and ER show decreasing trends, among which the anti-correlation between the ER of sample clusters and the number of their members is more significant. This indicates that the members’ number of the sample clusters has a significant impact on their ER, and the sample clusters with a large number of members have much less elliptical internal morphology than external morphology. The result of a shrinking of the MD of sample clusters with the number of their members can be interpreted as the influence of self-gravitational binding of sample clusters on their morphologies, because the larger the number of member stars in a sample cluster is, the stronger its self-gravitational binding capacity is.
@@
Moreover, there may be a compact internal structure in a cluster with a large number of members. This structure is prone to smaller dynamical timescales, thus speeding up clusters’ evolution (including evaporation rates and the process of mass segregation) (Spitzer, 1969; Madrid et al., 2012). It is well-known that member stars in the evaporation process escape almost from the external region within the tidal radius of clusters through Lagrangian points of the Jacobi potential sphere [(e.g., Portegies Zwart et al., 2010, see their Fig 6) and (e.g., Krumholz et al., 2019, see their Fig 15)]. The Jacobi potential sphere of the cluster with a number of members that looks like an egg may be fully filled by its members, while its internal morphology tends to be compact and spherical due to two-body interaction (Chen et al., 2004). The analysis above suggests that the number of the member stars of sample clusters is most likely an important factor in maintaining their morphological coherence, implying that the gravitational binding of the clusters is responsible for keeping their structure stable and less susceptible to external perturbations.
@@
In addition, we explored whether age affects the morphological coherence of the sample clusters. As open clusters age, their member stars redistribute as a result of the two-body relaxation process, leading to morphological changes in their internal structures. It is theoretically speculated that the MD of the sample clusters varies with their ages. Older clusters are still alive. Apart from the disturbance of the external environment, the morphological stability of the clusters is most likely a talisman for their stable survival. We thus explore the relationship between the MD and ER of the sample clusters and their ages, as shown in the (b) subplot of Fig. 7.
@@
From the left panel of the (b) subplot of Fig. 7, we can see that as the age of sample clusters increases, their MD presents only a slight downward trend, or even no change if you look at fluctuations of the median filter line (red). This demonstrates that the MD of the sample clusters is independent of their age, which is inconsistent with our speculation above. This is because the age is only a time scale of clusters existing and evolving. Many clusters may dissolve at any age due to weaker self-gravitational bindings, stronger external perturbations, or collisions with giant molecular clouds. In spite of the external influence on the evolution of the sample clusters, the number of their members is a key factor for their existence. For different clusters with different numbers of members, they may have the same age, so we can find their different MDs at the same age.
@@
Figure 7: The morphological coherence (MD and ER) vs. the fundamental parameters. The open circles in all panels represent the sample clusters, with the black bars being the errors of the morphological coherence (MD and ER). The red curves indicate the median filter lines of all distributions.
@@ 3.3 Correlations of the morphological coherence of sample clusters with spatial position
Although the MD and ER of the sample clusters do not show any interesting distribution trends on the XY plane, it is worth exploring their distribution on the X-axis and Y-axis separately, because the morphology of the sample clusters may be affected differently at different spatial locations. Figure 9 shows the distributions of the MD and ER along the X-axis and Y-axis, respectively. We also plotted their median filter lines (red curves) in this picture. In the left panel of the (a) subplot of Fig. 9, we did not find any increasing or decreasing trend between the MD and the X-axis, which implies the MD of the sample clusters being independent of the X-axis. This result is not consistent with our expectations, because along the X-axis, the closer to the Galactic center, the more disturbed the morphology of the sample clusters are, hence the larger their MDs should be. Of course, it is likely due to a small number of sample clusters. Therefore, we need more open clusters to verify this result in the future, if possible. In addition, we also found that the MDs of the sample clusters are evenly distributed along the Y-axis. Since along the Y-axis, the distributions of the sample clusters are almost symmetric in the regions of its positive and negative semi-axes, the MD has nothing to do with the Y-axis.
Furthermore, the ER of the sample clusters presents fluctuations along the X-axis toward the Galactic center, as displayed in the left panel of the (b) subplot of Fig. 9, but it does not show any trend at a whole. This means that the ER of the sample clusters is sensitive to the external environment along the X-axis if sample effects are not considered. To say the least, the external morphology of the sample clusters changes along the X-axis due to changes in external gravitational perturbations. The closer one gets to the Galactocentric side, the stronger the potential in the external environment is, and the more perturbation the morphology of the sample clusters will suffer. Like the MDs of the sample clusters, their ER also does not have any significant relationship with the Y-axis, as shown in the right panel of the (b) subplot of Fig. 9.
Figure 8: Left: distribution of the sample clusters on the plane of the galactic disk. Each circle represents a sample cluster, with its size proportional to the MD of the sample, and its color is coded according to the members’ number of the sample clusters. The different colored dashed lines indicate the positions of the local arm detected by different types of tracers, the data involved was from Hao et al. (2021). Right: distribution of the sample clusters on the profile map of the disk. Each circle represents one cluster, with its size being proportional to the cluster’s MD, and its color being coded according to the cluster’s ER. The dot-dashed line denotes the mid-plane (Z $\geq$ -20.8 pc) of the disk.
@@
Figure 9: Morphological coherence (MD and ER) vs. Spatial position. All open circles (representing the sample clusters), all black error bars (being the errors of the MD or ER), and red curves (denoting median filter lines), same as those in Fig. 7. ‘GC’ and ‘DR’ represent the Galactic center and the Galactic differential rotation, respectively.
@@ 3.4 Correlations of the MD of sample clusters with external environment
Figure 10: Morphological coherence vs. External environmental parameters. All open circles (representing sample clusters), all black error bars (being errors of the MD or ER), red curves (indicating median filter lines), same as those in Fig. 7.
@@
In Subsection 3.3, although the MD of the sample clusters is not related to their spatial positions, the external environment around the positions may be different. Therefore, some forces embedded in the external environment may impact their morphological coherence. To explore it, we employ the Python module galpy (Bovy, 2015) to derive some parameters involving the external environment, based on the galpy.potential module, MWPotential2014. The axisymmetric potential module, derived by a simple model fit to the available dynamical data for our galaxy, comprises a Miyamoto-Nagai disk (Miyamoto & Nagai, 1975), a bulge, and a dark matter halo modeled with a Navarro-Frenk-White (NFW) potential (Navarro et al., 1997). Inputting the coordinate, proper motion, and radial velocity data of the sample clusters, and then we can obtain orbit eccentricities, radial forces (FR), and vertical forces (Fz) in cylindrical coordinates [^1] It is a cylindrical coordinate centered on Galactic center, following left-handed frame in galpy. of these sample clusters. The meaning of these parameters can be also found on the web [^2] https://docs.galpy.org/en/v1.7.0/.
@@
Since the sample clusters with large orbital eccentricities necessarily move across different potential energy surfaces in the galactic disk compared to those with small eccentricities, their MD and ER are consequentially affected to a greater extent. In the left panels of the (a) and (b) subplots of Fig. 10, inconsistent with our inference, we found that the orbital eccentricity of the sample clusters does not have any significant relationship with their MDs. It is also not significantly related to the ER of the sample clusters either, with only local ups and downs in the distribution between the ERs and the orbital eccentricities. This manifests that the ER may be more sensitive than the MD to the orbital eccentricities of the sample clusters. Therefore, as the sample clusters move across the Galactic disk, their crossing of different potential energy surfaces may affect their ER, but not linear (see red curve in the left panels of the (b) subplot of Fig. 10).
@@
The median panels of the (a) and (b) subplots of Fig. 10 show the distributions between the radial forces (FR) and the MDs of the sample clusters and their ERs, respectively. We found that the MDs of the sample clusters keep no change as the FR increases, with their ER fluctuating. It implies that the radial forces can not impact the MDs of the sample clusters, but may cause their ERs to exhibit fluctuations by affecting their morphology. Here we still keep sample effects in our minds. Therefore, this result needs to be tested by a larger sample in the future.
@@
Moreover, apart from the radial forces, we also explored whether vertical forces are related to the MDs and ERs of the sample clusters. However, we only detected a fluctuation of the distribution between the ERs of the sample clusters and their vertical forces (Fz), with no apparent relationship among them, as shown in the right panel of (b) subplot of Fig. 10. The vertical forces are independent of the MD of the sample clusters, as shown in the right panel of (a) subplot of Fig. 10, which suggests that the MD is not affected by the vertical forces. In addition, the fluctuation between the ER and the Fz may be due to the insufficient sample, so we need to treat this fluctuation with caution.

### v2

@@ 3.1 Distribution of morphological coherence of sample clusters
We obtained all parameters, including the MD and ER of 132 sample clusters, through ellipsoid fitting and expressions above defined, compiled a machine-readable format and seen in Appendix 1. Figure 4 displays the distributions of the MD with its error (left panel) as well as the ER with its error (right panel). We found that the errors of the MD of most sample clusters are less than 0.1, with most ER errors less than 0.2. Meanwhile, both errors, expressed as $\sigma_{d}$ and $\sigma_{e_{core}/e_{outer}}$, respectively, present relatively apparent positive correlations with the MD and ER. The main reason is that the fitting error is larger for the sample clusters with large MD or large ER.
@@
To get a preliminary view of the morphological coherence distribution of the sample clusters, we plotted the distribution between the MD and the ER of the sample clusters, as shown in Fig. 5. Each colored circle represents one sample cluster, with its color representative of the number of the sample clusters’ members. The gray bars (horizontal and vertical directions) cross it are the error bars of MD and ER, respectively. The horizontal dashed line in the Figure marks the position of ER equal to 1, where the ellipticity of the cluster’s internal morphology is consistent with the ellipticity of its external morphology. The picture shows an increasing tendency of the ER ($e_{core}$/$e_{outer}$) values as a function of the MD ($d$), which implies that sample clusters with more elliptical external morphology than internal morphology typically present smaller morphological dislocations (MD), as shown by the dark red-filled circles in Fig. 5; thus, they have greater morphological stability. Meanwhile, it is apparent that these clusters have a larger number of member stars. Therefore, we concluded that the number of sample clusters’ members plays a crucial role in keeping their morphological stability.
@@
Figure 5: Distribution of the MD of the sample clusters and their ER. Each colored circle represents one sample cluster, with its color being coded according to the logarithmic value of the number of member stars of the sample clusters. The horizontal and vertical gray bars indicate the errors in the MD of the clusters and their ER, respectively. The dashed line denotes the position of the smallest morphological difference between the clusters’ inner and outer regions.
@@
In addition, we found (as seen in Fig. 5) a sample cluster (HSC 2986) that departs from the distribution of most sample clusters. It is a special object for our sample population because it has a relatively small number of members and the largest MD among all samples, although the error of the cluster’s MD is large. To examine the cluster further, we traced its 3D spatial distribution, as shown in Fig. 6. The left panel of the Figure presents the 3D spatial distributions and the fitted ellipsoids of the internal and external structures of the HSC 2986. Its internal and external structures were divided based on the radius of the half number of cluster members. In this panel, the black dashed line indicates the position of the MD of HSC 2986, with its length representing, the unnormalized MD (see Eq. 1). To verify whether the cluster really has a large unnormalized MD, we adopted the Gaussian mixture model (GMM) used by Tarricq et al. (2022) to distinguish among open clusters’ core region, tidal tail, and corona in two dimensions, with 1000 iterations to divide the members’ distribution of HSC 2986 into two components, as shown in the right panel of Fig. 6. The two components are purple and black dots in this picture. We then evaluated the density centers (purple pentagram and black pentagram) in the two parts of the cluster separately, finally, we derived its unnormalized MD (see the red dashed line in the right panel of Fig. 6).
@@
Figure 6: 3D spatial structure of HSC 2986 and its ellipsoidal curves (left). The black dashed line indicates the unnormalized MD of HSC 2986. Other symbols and curves are the same as those in Fig. 2. 3D spatial structure of HSC 2986 (right). The purple and black dots are divided by the Gaussian mixture model (GMM) method, with the red dashed line meaning the unnormalized MD of HSC 2986. The purple and black pentagrams are the maximum points of the density of the purple and black scatter distribution.
@@
By comparing the unnormalized MDs of HSC 2986 obtained, we are able to determine that HSC 2986 has a large unnormalized MD, which implies that the cluster is weakly morphologically stable. Moreover, we can see that the whole members’ distribution of HSC 2986 does not show a distinct core, presenting instead a diffuse spatial arrangement. Therefore, we speculated that the cluster may be susceptible to disruption by external forces, regardless of the cluster’s evolutionary state. However, it is also possible that its core and outer structures are two separate groups due to the large unnormalized MD if HSC 2986 is misclassified as an open cluster.
@@
3.2 Correlations of the morphological coherence of sample clusters with the number of members and age
@@
Figure 7 displays the distributions between the morphological coherence (MD and ER) and the number of the clusters’ members as well as their ages, with their median filter curves. It is apparent (as seen in Fig. 7a) that as the number of clusters’ members increases, both their MD and ER show decreasing trends; among them, the anti-correlation between the ER of sample clusters and the number of their members is more significant. This indicates that the members’ number of the sample clusters has a significant impact on their ER and the sample clusters with a large number of members have much less elliptical internal morphology than external morphology. The result of a shrinking of the MD of sample clusters with the number of their members can be interpreted as the influence of self-gravitational binding of sample clusters on their morphologies, because the larger the number of member stars in a sample cluster, the stronger its self-gravitational binding capacity.
@@
Moreover, there may be a compact internal structure in a cluster with a large number of members. This structure is prone to smaller dynamical timescales, thereby speeding up the cluster evolution, including the evaporation rates and the process of mass segregation (Spitzer, 1969; Madrid et al., 2012). It is well known that member stars in the evaporation process escape almost from the external region within the tidal radius of clusters through Lagrangian points of the Jacobi potential sphere (e.g., Portegies Zwart et al., 2010; Krumholz et al., 2019, see respectively their Fig. 6 and Fig. 15). The Jacobi potential sphere of the cluster with a number of members that looks like an egg may be fully filled by its members, while its internal morphology tends to be compact and spherical due to two-body interaction (Chen et al., 2004). The analysis above suggests that the number of the member stars of sample clusters is most likely an important factor in maintaining their morphological coherence, implying that the gravitational binding of the clusters is responsible for keeping their structure stable and less susceptible to external perturbations.
@@
In addition, we explored whether age affects the morphological coherence of the sample clusters. As open clusters age, their member stars redistribute as a result of the two-body relaxation process, leading to morphological changes in their internal structures. It has been theoretically speculated that the MD of the sample clusters varies with their ages. Older clusters are still alive. Apart from the disturbance of the external environment, the morphological stability of the clusters is most likely a talisman for their stable survival. In this way, we explored the relationship between the MD and ER of the sample clusters and their ages, as shown in the (b) subplot of Fig. 7.
@@
From the left panel of the (b) subplot of Fig. 7, we can see that as the age of sample clusters increases, their MD presents only a slight downward trend, or even no change if we are looking at fluctuations of the median filter line (red). This demonstrates that the MD of the sample clusters is independent of their age, which is inconsistent with our speculation above. This is because the age is only a time scale of clusters existing and evolving. Many clusters may dissolve at any age due to weaker self-gravitational bindings, stronger external perturbations, or collisions with giant molecular clouds. In spite of the external influence on the evolution of the sample clusters, the number of their members is a key factor for their existence. For different clusters with different numbers of members, they may have the same age; thus, we can find their different MDs at the same age.
@@
Figure 7: Morphological coherence (MD and ER) versus the fundamental parameters. The open circles in all panels represent the sample clusters, with the black bars being the errors of the morphological coherence (MD and ER). The red curves indicate the median filter lines of all distributions.
@@ 3.3 Correlations of the morphological coherence of sample clusters with spatial position
Although the MD and ER of the sample clusters do not show any interesting distribution trends on the XY plane, it is worth exploring their distribution on the X-axis and Y-axis separately, because the morphology of the sample clusters may be affected differently at different spatial locations. Figure 9 shows the distributions of the MD and ER along the X-axis and Y-axis, respectively. We also plotted their median filter lines (red curves) in this picture. In the left panel of the (a) subplot of Fig. 9, we did not find any increasing or decreasing trend between the MD and the X-axis, which implies the MD of the sample clusters being independent of the X-axis. This result is not consistent with our expectations because along the X-axis, the closer to the Galactic center, the more disturbed the morphology of the sample clusters are; hence, the larger their MDs are expected to be. Of course, it is likely due to a small number of sample clusters. Therefore, we need more open clusters to verify this result in the future, if possible. In addition, we also found that the MDs of the sample clusters are evenly distributed along the Y-axis. Since along the Y-axis, the distributions of the sample clusters are almost symmetric in the regions of its positive and negative semi-axes, the MD has nothing to do with the Y-axis.
Furthermore, the ER of the sample clusters presents fluctuations along the X-axis toward the Galactic center, as displayed in the left panel of the (b) subplot of Fig. 9, but it does not show any trend at a whole. This means that the ER of the sample clusters is sensitive to the external environment along the X-axis if sample effects are not considered. To say the least, the external morphology of the sample clusters changes along the X-axis due to changes in external gravitational perturbations. The closer we get to the Galactocentric side, the stronger the potential in the external environment and the more perturbation the morphology of the sample clusters will suffer. Like the MDs of the sample clusters, their ER also does not have any significant relationship with the Y-axis, as shown in the right panel of the (b) subplot of Fig. 9.
Figure 8: Distribution of the sample clusters on the plane of the galactic disk (left). Each circle represents a sample cluster, with its size proportional to the MD of the sample, and its color is coded according to the members’ number of the sample clusters. The different colored dashed lines indicate the positions of the local arm detected by different types of tracers, the data involved was from Hao et al. (2021). Distribution of the sample clusters on the profile map of the disk (right). Each circle represents one cluster, with its size being proportional to the cluster’s MD and its color coded according to the cluster’s ER. The dot-dashed line denotes the mid-plane (Z $\geq$ -20.8 pc) of the disk.
@@
Figure 9: Morphological coherence (MD and ER) versus Spatial position. All open circles (representing the sample clusters), all black error bars (being the errors of the MD or ER), and red curves (denoting median filter lines), same as those in Fig. 7. “GC” and “DR” represent the Galactic center and the Galactic differential rotation, respectively.
@@ 3.4 Correlations of the MD of sample clusters with external environment
Figure 10: Morphological coherence versus external environmental parameters. All open circles (representing sample clusters), all black error bars (being errors of the MD or ER) and red curves (indicating median filter lines) are the same as those in Fig. 7.
@@
In Section 3.3, although the MD of the sample clusters is not related to their spatial positions, the external environment around the positions may be different. Therefore, some forces embedded in the external environment may impact their morphological coherence. To explore it, we employ the Python module galpy (Bovy, 2015) to derive some parameters involving the external environment, based on the galpy.potential module: MWPotential2014. The axisymmetric potential module, derived by a simple model fit to the available dynamical data for our galaxy, comprises a Miyamoto-Nagai disk (Miyamoto & Nagai, 1975), a bulge, and a dark matter halo modeled with a Navarro-Frenk-White (NFW) potential (Navarro et al., 1997). Inputting the coordinate, proper motion, and radial velocity data of the sample clusters, and then we can obtain orbit eccentricities, radial forces (FR), and vertical forces (Fz) in cylindrical coordinates [^1] It is a cylindrical coordinate centered on Galactic center, following left-handed frame in galpy. of these sample clusters. The meaning of these parameters can be also found on the web [^2] https://docs.galpy.org/en/v1.7.0/.
@@
Since the sample clusters with large orbital eccentricities necessarily move across different potential energy surfaces in the galactic disk compared to those with small eccentricities, their MD and ER are consequentially affected to a greater extent. In the left panels of the (a) and (b) subplots of Fig. 10, inconsistent with our inference, we found that the orbital eccentricity of the sample clusters does not have any significant relationship with their MDs. It is also not significantly related to the ER of the sample clusters either, with only local ups and downs in the distribution between the ERs and the orbital eccentricities. This manifests that the ER may be more sensitive than the MD to the orbital eccentricities of the sample clusters. Therefore, as the sample clusters move across the Galactic disk, their crossing of different potential energy surfaces may affect their ER, but not linear (see red curve in the left panels of Fig. 10b).
@@
The median panels of the (a) and (b) subplots of Fig. 10 show the distributions between the radial forces (FR) and the MDs of the sample clusters and their ERs, respectively. We found that the MDs of the sample clusters keep no change as the FR increases, with their ER fluctuating. It implies that the radial forces can not impact the MDs of the sample clusters, but may cause their ERs to exhibit fluctuations by affecting their morphology. Here, we still retain the sample effects in our consideration. Therefore, this result needs to be tested by a larger sample in the future.
@@
Moreover, apart from the radial forces, we also explored whether vertical forces are related to the MDs and ERs of the sample clusters. However, we only detected a fluctuation of the distribution between the ERs of the sample clusters and their vertical forces (Fz), with no apparent relationship among them, as shown in the right panel of Fig. 10b. The vertical forces are independent of the MD of the sample clusters, as shown in the right panel of Fig. 10a, which suggests that the MD is not affected by the vertical forces. In addition, the fluctuation between the ER and the Fz may be due to the insufficient sample, so we need to treat this fluctuation with caution.